OpenStudy (anonymous):

i need to find an equation of a parabola with the dimensions of the st louis arch. its 630ft high and 630ft wide.

OpenStudy (mathmale):

Draw your parabola. Label the maximum value of y. Identify the y-intercept of the parabola. Determine the value of x at which y = 0. Consider the equation y=ax^2 + bx + c OR y = c*x^2 + k. Which one would be the easier to use here?

OpenStudy (anonymous):

the first one

OpenStudy (jdoe0001):

\(\bf recall\implies y={\color{red}{ a}}(x-h)^2+k\qquad vertex\ (h,k)\)

OpenStudy (mathmale):

JDoe's illustration and his equation are both very helpful. Would you, @seesawn, please identify the vertex of the parabola you're trying to describe with an equation. Referring to JDoe's equation, above, what is h? what is k? Knowing the coordinates of the vertex, and the right x-intercept, how would you go about finding the value of the coefficient " a " in his equation?

OpenStudy (anonymous):

what does a equal?

OpenStudy (anonymous):

h=0 k=630

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

so what's "a" ? well, we dunno but we know from the points (-315, 0) and (315, 0) that when x = \(\pm 350\) y = 0 so we could just use either point to get "a" thus \(\bf y={\color{red}{ a}}(x-0)^2+630\qquad vertex\ (0,630) \\ \quad \\ x = -315\qquad y=0\implies y={\color{red}{ a}}(x-0)^2+630\implies 0={\color{red}{ a}}(630-0)^2+630\) solve for "a"

OpenStudy (jdoe0001):

ahemm rather when x = \(\pm 315\) y = 0 I meant =) anyhow, so you'd use either point to solve for "a"

OpenStudy (jdoe0001):

hmmm I even used the wrong haemm lemme fix all that,shoot, the typos

OpenStudy (jdoe0001):

\(\bf y={\color{red}{ a}}(x-0)^2+630\qquad vertex\ (0,630) \\ \quad \\ x = -315\qquad y=0\implies y={\color{red}{ a}}(x-0)^2+630\implies 0={\color{red}{ a}}(-315-0)^2+630\)

OpenStudy (jdoe0001):

shoot... ahemm ok \(\bf y={\color{red}{ a}}(x-0)^2+630\qquad vertex\ (0,630) \\ \quad \\ x = -315\qquad y=0\implies y={\color{red}{ a}}(x-0)^2+630 \\\quad \\\implies 0={\color{red}{ a}}(-315-0)^2+630\)

OpenStudy (jdoe0001):

so using the point (-315, 0) and the vertex values, solve for "a" :)

OpenStudy (anonymous):

would it be 157.5? or did i do that wrong

OpenStudy (jdoe0001):

well... ahemm.... sure?

OpenStudy (anonymous):

if i got it right, the equation would be y=157.5(-315-0)^2+630

OpenStudy (anonymous):

right?

OpenStudy (jdoe0001):

hmmm well.. ahemm 157.5?

OpenStudy (jdoe0001):

\(\bf y={\color{red}{ a}}(x-0)^2+630\qquad vertex\ (0,630) \\ \quad \\ x = -315\qquad y=0\implies y={\color{red}{ a}}(x-0)^2+630 \\\quad \\\implies 0={\color{red}{ a}}(-315-0)^2+630\) ... what would you get for "a"?

OpenStudy (anonymous):

idk, i subtracted the 630, then distributed the square, divided the answer from the parenthesis, and got 157.5

OpenStudy (jdoe0001):

hmmm the 630 is outside the parenthesis, recall your PEMDAS :)

OpenStudy (anonymous):

shiiiiizzzzzzzzz

OpenStudy (anonymous):

so wouldnt that give me 99855?

OpenStudy (jdoe0001):

yes, it would

OpenStudy (jdoe0001):

well, actaully no

OpenStudy (jdoe0001):

anyhow -> \(\bf x = -315\qquad y=0\implies y={\color{red}{ a}}(x-0)^2+630 \\\quad \\\implies 0={\color{red}{ a}}(-315-0)^2+630 \\ \quad \\ 0=a(-315)^2+630\implies -630=a(-315)^2\implies \cfrac{-630}{(-315)^2}=a\)

OpenStudy (anonymous):

-0.00634920634?

OpenStudy (jdoe0001):

then you can simplify that, and that would be "a" :) well, you may want to leave it as fraction, rather than decimal, either will work I gather

OpenStudy (anonymous):

ok thanks for the help!!!

OpenStudy (jdoe0001):

\(\bf 0=a(-315)^2+630\implies -630=a(-315)^2\implies \cfrac{-630}{(-315)^2}=a \\ \quad \\ -\cfrac{2}{315}=a\qquad thus \\ \quad \\ y={\color{red}{ a}}(x-0)^2+630\quad vertex\ (0,630)\implies y={\color{red}{ -\cfrac{2}{315}}}(x-0)^2+630\)

OpenStudy (jdoe0001):

yw